OpenStudy (anonymous):

The width of a rectangular mirror is 3/4 the measure of the length of the mirror. If the area is 192 in.squared, what are the length and width of the mirror?

OpenStudy (jack1):



OpenStudy (jack1):



OpenStudy (jack1):

Area of a rectangle = ...? @rwhovian98 ?

OpenStudy (anonymous):

Length times width?

OpenStudy (jack1):

sweet and your width is 3L/4 so times it out 192 = L x W 192 = L x 3L/4 ... u finish is up

OpenStudy (jack1):

it should finish with: L = _______

OpenStudy (anonymous):

What do you mean by times it out?

OpenStudy (jack1):

\[192 = L \times \frac{ 3L }{ 4 } \] \[192 = \frac{ 3LL }{ 4 } \] \[192 = \frac{ 3L^2 }{ 4 } \] \[192 \times 4 = \frac{ 3L^2 }{ 4 } \times 4 \] \[768 = 3L^2 \] ...

OpenStudy (jack1):

so just finish that off until you're left with L = something then use that to find width (W = 3L/4)

OpenStudy (anonymous):

I get it now, thanks.

OpenStudy (jack1):

all good dude yw